# No Need to Look Back: An Efficient and Scalable Approach for Temporal Network Representation Learning

### v1

@@ Abstract
Temporal graph representation learning (TGRL) is crucial for modeling complex, dynamic systems in real-world networks. Traditional TGRL methods, though effective, suffer from high computational demands and inference latency. This is mainly induced by their inefficient sampling of temporal neighbors by backtracking the interaction history of each node when making model inference. This paper introduces a novel efficient TGRL framework, No-Looking-Back (NLB). NLB employs a “forward recent sampling” strategy, which bypasses the need for backtracking historical interactions. This strategy is implemented using a GPU-executable size-constrained hash table for each node, recording down-sampled recent interactions, which enables rapid response to queries with minimal inference latency. The maintenance of this hash table is highly efficient, with $O(1)$ complexity. NLB is fully compatible with GPU processing, maximizing programmability, parallelism, and power efficiency. Empirical evaluations demonstrate that NLB matches or surpasses state-of-the-art methods in accuracy for link prediction and node classification across six real-world datasets. Significantly, it is 1.32-4.40 $\times$ faster in training, 1.2-7.94 $\times$ more energy efficient, and 1.97-5.02 $\times$ more effective in reducing inference latency compared to the most competitive baselines. The link to the code: https://github.com/Graph-COM/NLB.
Machine Learning, ICML
@@ 1 Introduction
Temporal networks are extensively employed to model real-world complex systems. Such networks dynamically evolve as interactions occur or states change within their nodes. Learning node representations in these networks is pivotal, as they can be used in many downstream tasks such as monitoring or predicting changes in the temporal networks. Representative applications include forecasting interactions between node pairs for the tasks such as anomalous transaction detection in financial networks (Ranshous et al., 2015; Wang et al., 2021a; Li et al., 2021), recommendations in social networks (Liben-Nowell & Kleinberg, 2007) and in e-commerce systems (Koren, 2009). Node representations can also be used to predict individual nodes’ properties such as community detection (Bhagat et al., 2011; Kazemi et al., 2019b) and fraudsters detection (Wang et al., 2021b; Li et al., 2017; Ding et al., 2019; Yuan et al., 2021)
The interaction patterns surrounding a node often provide key indicators of its present state. Inspired by Graph Neural Networks (GNNs) (Kipf & Welling, 2017; Hamilton et al., 2017) that may effectively encode the neighborhood of a node in static graphs, researchers have developed models for temporal networks to incorporate both the structural and temporal aspects of a node’s historical neighborhood. These methods can be collectively named as temporal graph representation learning (TGRL) (Rossi et al., 2020; Wang et al., 2021d; Xu et al., 2020; Luo & Li, 2022; Kumar et al., 2019; Trivedi et al., 2019; Wang et al., 2021c).
@@
We identify that for most state-of-the-art (SOTA) TGRL methods, the online collection of the information around a node’s neighborhood is a major computational bottleneck. As a query comes in, current methods have to backtrack in time, sample a subset of historical interactions around the node of interest, and aggregate the information from these down-sampled neighboring interactions for prediction. Sampling neighboring nodes is a common strategy in representation learning for large static graphs (Hamilton et al., 2017; Ying et al., 2018; Huang et al., 2018). However, this strategy faces challenges when applied to temporal networks, as they do not incorporate temporal information and may encounter high latency due to the enlarged sample space created by neighbors across various timestamps.
@@
TGAT (Xu et al., 2020)
TGN (Rossi et al., 2020)
CAWN (Wang et al., 2021d)
APAN (Wang et al., 2021c)
NAT (Luo & Li, 2022)
@@
Table 1: Comparisons between several recent TGRL methods.
To address this, a strategy termed recent sampling has been proposed (Xu et al., 2020; Wang et al., 2021d), where more recent interactions are given higher weights in sampling, such as through a probability proportional to $\exp(-c\triangle t)$, with $c\geq 0$ and $\triangle t$ reflecting the time difference, indicating the recency of an interaction. While recent sampling can improve prediction performance, the process of calculating sampling weights and executing non-uniform sampling has high computational complexity $O(|\mathcal{N}_{u}^{t}|)$ where $|\mathcal{N}_{u}^{t}|$ denotes the size of the historical neighbors of node $u$ till time $t$. This significantly slows down the system. Consequently, most TGRL methods (Rossi et al., 2020; Wang et al., 2021c; Zhou et al., 2022) opt for either simple uniform sampling or a strategy named truncation, where only a fixed number of most recent neighbors are considered, to reduce the complexity. Even with such simplification, the inference latency may still be high because backtracking and sampling of these historical interactions are performed within CPUs instead of GPUs due to the inherent irregularity. A recent framework called TGL (Zhou et al., 2022) reduced the issue via leveraging powerful multi-core CPUs complemented by meticulously crafted C++ parallelizable programming for fine-grained memory and thread management. However, multi-threading in CPUs is less power-efficient and may still suffer from longer latency than parallelism in GPUs.
@@
NLB supports two variants NLB-edge and NLB-node, which differ in the used hash keys being either link ID or node ID. NLB-edge and NLB-node provably achieve recent sampling of links and recent sampling of nodes respectively while both support node-level and link-level applications. Empirically, NLB is comparable or improves SOTA link prediction and node classification accuracy in 6 real-world datasets while being 1.32-4.40$\times$ faster in training and 1.2-7.94$\times$ more energy efficient than the three most competitive baselines, and 1.97-5.02$\times$ more efficient in inference latency than the fastest baseline that leverages powerful multi-thread CPUs to perform backward sampling.
@@ 2 More Related Works
There are two streams of work on representation learning for temporal networks. Earlier work proposed to aggregate the sequence of links into network snapshots and processing a sequence of static network snapshots (Zhou et al., 2018; Du et al., 2018; Mahdavi et al., 2018; Singer et al., 2019; Nguyen et al., 2018; Sankar et al., 2020), but they may suffer from low prediction accuracy as they cannot capture time-sensitive information within static snapshots.
More recent temporal graph learning methods process link streams directly (Trivedi et al., 2017, 2019; Kumar et al., 2019; Souza et al., 2022). Most of these methods learn representation for a node using the historical information from the neighboring nodes it previously interacted with (Xu et al., 2020; Rossi et al., 2020; Wang et al., 2021d; Zhou et al., 2022; Luo & Li, 2022).
Due to the high computation overhead of temporal graph learning as reviewed in Sec.1, several recent works proposed improvements from the scalability perspective (Wang et al., 2021c; Zhou et al., 2022; Luo & Li, 2022). APAN (Wang et al., 2021c) proposed asynchronous graph propagation which passes information to most recent temporal neighbors and works with a sampling property similar to truncation without sampling. It improves inference latency but requires significant memory for buffering messages and the recent truncation strategy hurts prediction accuracy. A recent work NAT (Luo & Li, 2022) aims to improve link prediction performance and scalability by learning neighborhood-aware representations, which elegantly avoids backward sampling. However, its sampling property is not mathematically explained, and it relies on constructing joint neighborhood structural features of node pairs, which cannot be generalized to node level tasks. NAT inspires our sampling strategy NLB-node, while our framework can be applied to node-level tasks and support another strategy NLB-edge that often yields better performance than NLB-node.
@@ 4.1 The sampling strategies of existing works
While different approaches have different implementations, a common computational bottleneck for generating node representations during both training and inference is the online sampling of temporal neighbors. Existing sampling strategies consist of two steps: (1) Backtrack and collect the historical interactions of a node from the current timestamp; (2) Sample a subset of temporal neighbors from those interactions. Traditional methods (Xu et al., 2020; Rossi et al., 2020; Wang et al., 2021d; Zhou et al., 2022) perform these operations within CPUs, causing inefficiency in both time and power consumption for both training and inference.
To give an example, to learn the representations for node $u$ and $v$ at $t_{4}$ for downstream predictions (Fig. 1 Left), traditional methods perform backtracking and sampling sampling within the CPUs (Fig. 1 Middle). These blocking operations not only slow down the training and inference processes, but also introduce costly energy consumption from both CPU computation and CPU-GPU communication. Previous methods commonly adopted the following sampling strategies.
@@
Truncation of a fixed number of most recent neighbors, or simply named truncation, is a common strategy used by TGN (Rossi et al., 2020), APAN (Wang et al., 2021c) and TGL (Zhou et al., 2022). It does not sample with randomness but truncates the most recent $s$ temporal neighbors that appear in the history instead. Truncation is effective when the most recent neighbors provide sufficient information about the current state of a node’s neighborhood. However, as it does not allow randomness, it may fail to capture information that appears earlier in the history. In an extreme, if all most recent interactions are from the same neighbor, truncation neglects information from all other early neighbors.
@@
Uniform sampling is also commonly supported by TGN, TGL, TGAT (Xu et al., 2020), etc. It samples among all temporal neighbors in the past with equal probabilities. It is effective in capturing early-day neighbors which may benefit prediction. However, for many real-world applications, more recent events typically play a more important role. Moreover, uniform sampling, albeit straightforward, still requires the access to all of the historical temporal interactions. Since different nodes have different numbers of historical interactions, uniform sampling can hardly be implemented in the GPUs for parellelism, and thus is slower than truncation.
@@ 4.2 Our approach: Forward recent sampling
Recent sampling assigns larger sampling weights to more recent temporal neighbors. The probability a temporal neighbor of node $u$, e.g., $(v,e^{t^{\prime}}_{u,v},t^{\prime})$ is sampled for learning $Z_{u}^{t}$ is proportional to $\exp(c(t^{\prime}-t))$ where $c\geq 0$ is a constant. This sampling strategy is previously considered in Xu et al. (2020) and Wang et al. (2021d). Recent sampling has shown to be effective for many scenarios as both recent and some early-day neighbors could provide information. By tuning $c$, recent sampling may implement uniform sampling ($c\rightarrow 0$) and truncation ($c\rightarrow\infty$). However, it is hard to implement efficiently via the traditional backward sampling. Not only the existing way to implement recent sampling needs to access all the timestamps of historical interactions to compute those non-uniform sampling probabilities, but also performing non-uniform sampling is rather irregular, which could be much slower than truncation and uniform sampling.
Forward Recent Sampling. For both uniform sampling and recent sampling, if they are achieved through backward sampling as implemented by current methods, the time complexity is at least $O(|N_{u}^{t}|)$ as they need to access the entire historical neighbors. We now introduce forward sampling that can implement recent sampling in $O(1)$. An overview of the algorithm is demonstrated in Fig. 1 Right.
@@
Formally, we denote $S_{u}$ to be the down-sampled temporal neighbors of node $u$ stored within a GPU hash table. For efficient lookup and update in batches, we allocate fixed-size GPU memory for these tables, where the size $s$, as a user-specified parameter, is the maximum number of temporal neighbors to be sampled. The current snapshot of table $S_{u}$ at a timestamp $t$ is denoted as $S_{u}^{t}$. $S_{u}$ consists of key-value pairs where the keys are unique. Each key-value pair can correspond to one of the temporal neighbors of $u$, e.g., $(v,e^{t^{\prime}}_{u,v},t^{\prime})\in\mathcal{N}_{u}^{t}$. While we store the edge features $e^{t^{\prime}}_{u,v}$ as the values, the choice of the keys is flexible: It can either be the pair of the neighbor ID and timestamp, i.e.,  $(v,t^{\prime})$, or just the neighbor ID $v$. We name the method using $(v,t^{\prime})$ as NLB-edge and the one using $v$ as NLB-node. Different choices of the keys have different sampling properties which we will discuss in Sec. 4.3.
@@
Example. In Fig. 1, when link $(u,v,t_{4})$ arrives, we initialize an update to $S_{u}$ and $S_{v}$, creating new snapshots $S_{u}^{t_{4}}$ and $S_{u}^{t_{4}}$. Specifically, for NLB-edge which uses tuples of neighbor ID and timestamp as keys, the new temporal neighbor $(v,e^{t_{4}}_{u,v},t_{4})$ of $u$ will be assigned to position
@@ 4.3 The properties of forward recent sampling
Suppose links come in for any node (e.g. $u$) by following a Poisson point process with a constant intensity (e.g. $\lambda$), and suppose a temporal neighbor $(v,e_{u,v}^{t_{i}},t_{i})$ is inserted into $S_{u}$ at time $t_{i}$, then $\Pr((v,e_{u,v}^{t_{i}},t_{i})\in S_{u}^{t})=\exp(\frac{\alpha\lambda}{s}(t_{i}-t))$. [^1] Poisson point process is a commonly used assumption to model communication networks Lavenberg (1983), and is also an assumption used by CAWN (Wang et al., 2021d). Proof. See Appendix A.
This result matches the definition of recent sampling (Def. 4.3) where the probability a temporal neighbor gets sampled is proportional to $\exp(c(t_{i}-t))$. If we instead suppose that each unique neighboring node (e.g. $v$) of $u$ connects with $u$ following a different Poisson point process (e.g. $\lambda_{v}$), then we recover the sampling probabilities of NLB-node.

### v2

# Scalable and Efficient Temporal Graph Representation Learning via Forward Recent Sampling

@@ Abstract
Temporal graph representation learning (TGRL) is essential for modeling dynamic systems in real-world networks. However, traditional TGRL methods, despite their effectiveness, often face significant computational challenges and inference delays due to the inefficient sampling of temporal neighbors. Conventional sampling methods typically involve backtracking through the interaction history of each node. In this paper, we propose a novel TGRL framework, No-Looking-Back (NLB), which overcomes these challenges by introducing a forward recent sampling strategy. This strategy eliminates the need to backtrack through historical interactions by utilizing a GPU-executable, size-constrained hash table for each node. The hash table records a down-sampled set of recent interactions, enabling rapid query responses with minimal inference latency. The maintenance of this hash table is highly efficient, operating with $O(1)$ complexity. Fully compatible with GPU processing, NLB maximizes programmability, parallelism, and power efficiency. Empirical evaluations demonstrate that NLB not only matches or surpasses state-of-the-art methods in accuracy for tasks like link prediction and node classification across six real-world datasets but also achieves 1.32-4.40$\times$ faster training, 1.2-7.94$\times$ greater energy efficiency, and 1.63-12.95$\times$ lower inference latency compared to competitive baselines.
The link to the code: https://github.com/Graph-COM/NLB.
@@ 1 Introduction
Temporal networks are extensively employed to model real-world complex systems. Such networks dynamically evolve as interactions occur or states change within their nodes. Learning node representations in these networks is pivotal, as they can be used in many downstream tasks such as monitoring or predicting changes in the temporal networks. Representative applications include forecasting interactions between node pairs for the tasks such as anomalous transaction detection in financial networks [1, 2, 3], recommendations in social networks [4] and in e-commerce systems [5]. Node representations can also be used to predict individual nodes’ properties such as community detection [6, 7, 8] and fraudsters detection [9, 10, 11, 12].
The interaction patterns surrounding a node often provide key indicators of its present state. Inspired by Graph Neural Networks (GNNs) [13, 14] that may effectively encode the neighborhood of a node in static graphs, researchers have developed models for temporal networks to incorporate both the structural and temporal aspects of a node’s historical neighborhood. These methods can be collectively named as temporal graph representation learning (TGRL) [15, 16, 17, 18, 19, 20, 21].
@@
We identify that for most state-of-the-art (SOTA) TGRL methods, the online collection of the information around a node’s neighborhood is a major computational bottleneck. As a query comes in, current methods have to backtrack in time, sample a subset of historical interactions around the node of interest, and aggregate the information from these down-sampled neighboring interactions for prediction. Sampling neighboring nodes is a common strategy in representation learning for large static graphs [14, 22, 23, 24]. However, this strategy faces challenges when applied to temporal networks, as they do not incorporate temporal information and may encounter high latency due to the enlarged sample space created by neighbors across various timestamps.
@@
TGAT [17]
TGN [15]
CAWN [16]
APAN [21]
NAT [18]
@@
Table 1: Comparisons between several TGRL methods.
To address this, a strategy termed recent sampling has been proposed [17, 16], where more recent interactions are given higher weights in sampling, such as through a probability proportional to $\exp(-c\triangle t)$, with $c\geq 0$ and $\triangle t$ reflecting the time difference, indicating the recency of an interaction. While recent sampling can improve prediction performance, the process of calculating sampling weights and executing non-uniform sampling has high computational complexity $O(|\mathcal{N}_{u}^{t}|)$ where $|\mathcal{N}_{u}^{t}|$ denotes the size of the historical neighbors of node $u$ till time $t$. This significantly slows down the system. Consequently, most TGRL methods [15, 21, 25] opt for either simple uniform sampling or a strategy named truncation, where only a fixed number of most recent neighbors are considered, to reduce the complexity. Even with such simplification, the inference latency may still be high because backtracking and sampling of these historical interactions are performed within CPUs instead of GPUs due to the inherent irregularity. A recent framework called TGL [25] reduced the issue via leveraging powerful multi-core CPUs complemented by meticulously crafted C++ parallelizable programming for fine-grained memory and thread management. However, multi-threading in CPUs is less power-efficient and may still suffer from longer latency than parallelism in GPUs.
@@
NLB supports two variants NLB-edge and NLB-node, which differ in the used hash keys being either link ID or node ID. NLB-edge and NLB-node provably achieve recent sampling of links and recent sampling of nodes respectively while both support node-level and link-level applications. Empirically, NLB is comparable or improves SOTA link prediction and node classification accuracy in 6 real-world datasets while being 1.32-4.40$\times$ faster in training and 1.2-7.94$\times$ more energy efficient than the three most competitive baselines, and 1.63-12.95$\times$ more efficient in inference latency than the fastest baseline that leverages powerful multi-thread CPUs to perform backward sampling.
@@ 2 More Related Works
There are two streams of work on representation learning for temporal networks. Earlier work proposed aggregating the sequence of links into network snapshots and processing a sequence of static network snapshots [26, 27, 28, 29, 30, 31], but they may suffer from low prediction accuracy as they cannot capture time-sensitive information within static snapshots. More recent temporal graph learning methods process link streams directly [32, 20, 19, 33, 34, 35]. Most of these methods learn representation for a node using the historical information from the neighboring nodes it previously interacted with [17, 15, 16, 25, 18, 36, 37, 38].
There are many works for graph learning that consider improvement on scalability and inference efficiency. For GNNs on static graphs, a stream of work focuses on efficient neighbor sampling [14, 22, 23, 24] which speeds up both training and inference, while some other works studies the acceleration for inference via knowledge distillation [39, 40, 41] or channel pruning [42]. However, these techniques cannot be easily generalized to improve the scalability for temporal networks. First, as temporal links come in as a sequence, it is non-scalable to construct and maintain the whole graph for every new link. Furthermore, these methods do not resolve the bottleneck of historical neighbor sampling on temporal graphs, but neighborhood information is crucial for graph representation learning.
For temporal graphs, due to the more significant computation overhead of TGRLs as reviewed in Sec. 1, several recent works focusing on acceleration and efficiency inference are proposed [21, 25, 18, 43]. APAN [21] proposed asynchronous graph propagation which passes information to the most recent temporal neighbors and works with a sampling property similar to truncation without sampling. It improves inference latency but requires significant memory for buffering messages and the recent truncation strategy may hurt prediction accuracy. A recent work NAT [18] aims to improve link prediction performance and scalability by learning neighborhood-aware representations, which elegantly avoids backward sampling. However, its sampling property is not mathematically explained, and it relies on constructing joint neighborhood structural features of node pairs, which cannot be generalized to node-level tasks. In experiment (Sec. 5), we find that by removing the joint neighborhood features from NAT, it can be applied to node-level tasks. But its node-level prediction performance is worse than NLB and it is still slow compared to NLB.
Finally, temporal graph benchmark (TGB) [44] is proposed to standardize the evaluation pipelines for TGRL. We build a separate evaluation pipeline outside of TGB because our focus involves computational evaluations on billion-scale datasets which has not been supported by TGB. We provide supplementary evaluation of NLB on TGB in Appendix F.
@@ 4.1 The sampling strategies of existing works
While different approaches have different implementations, a common computational bottleneck for generating node representations during both training and inference is the online sampling of temporal neighbors. Existing sampling strategies consist of two steps: (1) Backtrack and collect the historical interactions of a node from the current timestamp; (2) Sample a subset of temporal neighbors from those interactions. Traditional methods [17, 15, 16, 25] perform these operations within CPUs, causing inefficiency in both time and power consumption for both training and inference.
To give an example, to learn the representations for node $u$ and $v$ at $t_{4}$ for downstream predictions (Fig. 1 Left), traditional methods perform backtracking and sampling within the CPUs (Fig. 1 Middle). These blocking operations not only slow down the training and inference processes, but also introduce costly energy consumption from both CPU computation and CPU-GPU communication. Previous methods commonly adopted the following sampling strategies.
@@
Truncation of a fixed number of most recent neighbors, or simply named truncation, is a common strategy used by TGN [15], APAN [21] and TGL [25]. It does not sample with randomness but truncates the most recent $s$ temporal neighbors that appear in the history instead. Truncation is effective when the most recent neighbors provide sufficient information about the current state of a node’s neighborhood. However, as it does not allow randomness, it may fail to capture information that appears earlier in the history. In an extreme, if all most recent interactions are from the same neighbor, truncation neglects information from all other early neighbors.
@@
Uniform sampling is also commonly supported by TGN, TGL, TGAT [17], etc. It samples among all temporal neighbors in the past with equal probabilities. It is effective in capturing early-day neighbors which may benefit prediction. However, for many real-world applications, more recent events typically play a more important role. Moreover, uniform sampling, albeit straightforward, still requires the access to all of the historical temporal interactions. Since different nodes have different numbers of historical interactions, uniform sampling can hardly be implemented in the GPUs for parallelism, and thus is slower than truncation.
@@ 4.2 Our approach: Forward recent sampling
Recent sampling assigns larger sampling weights to more recent temporal neighbors. The probability a temporal neighbor of node $u$, e.g., $(v,e^{t^{\prime}}_{u,v},t^{\prime})$ is sampled for learning $Z_{u}^{t}$ is proportional to $\exp(c(t^{\prime}-t))$ where $c\geq 0$ is a constant. This sampling strategy is previously considered in Xu et al. [17] and Wang et al. [16]. Recent sampling has shown to be effective for many scenarios as both recent and some early-day neighbors could provide information. By tuning $c$, recent sampling may approximate uniform sampling (for a small $c$) and truncation (for a large $c$). However, it is hard to implement efficiently via the traditional backward sampling. Not only the existing way to implement recent sampling needs to access all the timestamps of historical interactions to compute those non-uniform sampling probabilities, but also performing non-uniform sampling is rather irregular, which could be much slower than truncation and uniform sampling.
Forward Recent Sampling. For both uniform sampling and recent sampling, if they are achieved through backward sampling as practiced by current methods, the time complexity is at least $O(|N_{u}^{t}|)$ as they need to access the entire historical neighbors. We now introduce forward sampling that implements recent sampling in $O(1)$. An overview of the algorithm is demonstrated in Fig. 1 Right.
@@
Formally, we denote $S_{u}$ to be the down-sampled temporal neighbors of node $u$ stored within a GPU hash table. For efficient lookup and update in batches, we allocate fixed-size GPU memory for these tables, where the size $s$, as a user-specified parameter, is the maximum number of temporal neighbors to be sampled. The current snapshot of table $S_{u}$ at a timestamp $t$ is denoted as $S_{u}^{t}$. $S_{u}$ consists of key-value pairs with unique keys. Each key-value pair corresponds to one of the temporal neighbors of $u$, e.g., $(v,e^{t^{\prime}}_{u,v},t^{\prime})\in\mathcal{N}_{u}^{t}$. While we store the edge features $e^{t^{\prime}}_{u,v}$ as the values, the choice of the keys is flexible: It can either be the pair of the neighbor ID and timestamp, i.e.,  $(v,t^{\prime})$, or just the neighbor ID $v$. We name the method using $(v,t^{\prime})$ as NLB-edge and the one using $v$ as NLB-node. Different choices of the keys have different sampling properties which we will discuss in Sec. 4.3.
@@
Example. In Fig. 1, when link $(u,v,t_{4})$ arrives, we initialize an update to $S_{u}$ and $S_{v}$, creating new snapshots $S_{u}^{t_{4}}$ and $S_{v}^{t_{4}}$. Specifically, for NLB-edge which uses tuples of neighbor ID and timestamp as keys, the new temporal neighbor $(v,e^{t_{4}}_{u,v},t_{4})$ of $u$ will be assigned to position
@@ 4.3 The properties of forward recent sampling
Suppose links come in for any node (e.g. $u$) by following a Poisson point process with a constant intensity (e.g. $\lambda$), and suppose a temporal neighbor $(v,e_{u,v}^{t_{i}},t_{i})$ is inserted into $S_{u}$ at time $t_{i}$, then $\Pr((v,e_{u,v}^{t_{i}},t_{i})\in S_{u}^{t})=\exp(\frac{\alpha\lambda}{s}(t_{i}-t))$. [^1] Poisson point process is a commonly used assumption to model communication networks Lavenberg [45], and is also an assumption used by CAWN [16].
This result matches the definition of recent sampling (Def. 4.3) where the probability a temporal neighbor gets sampled is proportional to $\exp(c(t_{i}-t))$. If we instead suppose that each unique neighboring node (e.g. $v$) of $u$ connects with $u$ following a different Poisson point process (e.g. $\lambda_{v}$), we recover the sampling probabilities of NLB-node.